Given a standard normal distribution with mean =136 and standard deviation = 39, find P(X>84)


1
Expert's answer
2022-01-25T14:24:16-0500

"X\sim N(136, 39^2)"


"P(X>84)=1-P(X\leq 84)"

"=1-P(Z\leq \dfrac{84-136}{39})=1-P(Z\leq-\dfrac{4}{3})"

"\approx1-P(Z\leq -1.3333)\approx0.9088"